OpenStudy (anonymous):

pulled it out of my head... don't even know why, but will see if I can solve it.

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \ln(\ln x)~dx\]\[\int\limits_{ }^{ } \ln(\ln x) dx=x \ln(\ln x)- \int\limits_{ }^{ }\frac{1}{\ln x} \frac{1}{x} x~dx\]

OpenStudy (anonymous):

you can see what I am doing,..

OpenStudy (anonymous):

\[v=\ln(x)~~~~x=e^v,~~~~dv=\frac{1}{x}~dx\]

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \ln(\ln x)~dx=x \ln(\ln x) - \int\limits_{ }^{ }\frac{e^v}{v}~dv\]

hartnn (hartnn):

\(\Large \int \dfrac{1}{\ln x} dx\) there is no answer to this integral in terms of standard functions

OpenStudy (anonymous):

ohh, I did this before, it came out before in previous questions but it was just e^x/x dx, \[\int\limits_{ }^{ } \ln(\ln x)~dx=x \ln(\ln x) - \sum_{n=0}^{\infty} \frac{n!}{v^n}\]

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \ln(\ln x)~dx=x \ln(\ln x) - \sum_{n=0}^{\infty} \frac{n!}{(\ln x)^n}~\]

OpenStudy (anonymous):

I have dealt 2 problems before with the e^x/x.

hartnn (hartnn):

ofcourse you can use the series :) i was talking about standard functions like log f(x), e^{f(x)} , sin / cos /tan f(x), polynomials

OpenStudy (anonymous):

oh no, series only.... good that x came to be e^v.

OpenStudy (anonymous):

now though, I remember the integral of e^x/x .